OpenStudy (anonymous):

Please Help!! Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 2i

OpenStudy (anonymous):

one part will look like \[(x-5)(x+3)\]

OpenStudy (anonymous):

there are a few ways to find a quadratic with zeros \(-1+2i\) and \(-1-2i\) one way is to work backwards

OpenStudy (anonymous):

put \[x=-1+2i\] add 1\[x+1=2i\] square both sides (carefully)\[(x+1)^2=(2i)^2\\x^2+2x+1=-4\] add 4 \[x^2+2x+5\]

OpenStudy (anonymous):

your answer will be what you get when you multiply \[(x-5)(x+3)(x^2+2x+5)\]

OpenStudy (anonymous):

ok thank you!!!!

OpenStudy (anonymous):

yw btw i would cheat for the last part http://www.wolframalpha.com/input/?i=%28x-5%29%28x%2B3%29%28x^2%2B2x%2B5%29

OpenStudy (anonymous):

haha thats exactly what i did XD